7(b) Let X X Xn

, , , 1 2 K be random sample of size n from a distribution with probability 

density function 

θ < < θ >

=

θ−

,0 else where.

0, ,1 0

( )0,;

1 X X

f X Obtain a maximum likeyhood 

estimator of θ.


1
Expert's answer
2021-03-30T06:27:27-0400

Given probability distribution function is-


"f(x;θ)=θ^{−1}_2e^{−\dfrac{(x−θ_1)}{θ_2}}"


. The likelihood function is the density function regarded as a function of θ.


"L(\u03b8|x) = f(x|\u03b8), \u03b8 \u2208 \u0398. (1)"


The maximum likelihood estimator (MLE),


"\hat{θ}(x) = arg maxL(θ|x)."


for "x>\u03b8_1,\u03b8_2>0" Likelihood:


"L(θ)=L(θ∣X_1,X_2,…,X_n)=∏_{i=1}^n\dfrac{e^{\dfrac{−(x_i−θ_1)}{θ_2}}}{θ_2}"

factored out "\u03b8_2" as well as turned to product into a summation to get:


"=\dfrac{1}{θ^n_2}e^{∑^n_{i=1}}\dfrac{−(x_i−θ_1)}{θ_2}"


"ℓ(θ)=ln(L(θ))=\dfrac{1}{θ^n_2}e^{∑^n_{i=1}}\dfrac{−(x_i−θ_1)}{θ_2}"


"=ln(θ^{−n}_2)+∑_{i=1}^n−\dfrac{(x_i−θ_1)}{θ_2}"


   "=−nln(θ_2)−\dfrac{∑^n_{i=1}x_i−nθ_1}{θ_2}"


    "=−nln(θ_2)−\dfrac{∑^n_{i=1}x_i}{θ_2}+\dfrac{nθ_1}{\theta_2}"